# Revised Mass and Orbit of Eridani b: A 1 Jupiter-Mass Planet on a Near-Circular Orbit

### v1

@@ 4.3.2 The GP results are robust to the choice of data
While our main models used separate GPs for each instrument, we also tested a model with a single GP for all data. This setup, related to a strategy suggested in Blunt etĀ al. (2023), shares temporal correlation information between instruments, instead of treating data from each instrument as uncorrelated. We find essentially the same low-eccentricity results from this model, indicating that the GPs do not appear to be overfitting the data from each instrument.
6em
@@ 4.3.3 High cadence data are the source of the disagreement
-1.25cm-0.5cm
The HARPS (pre) series contains a burst of high cadence data, with 430 out of all 530 points captured over a 2-night range in October, 2004. This dense data from HARPS visibly traces out a correlated signal with a timescale much higher than the Keplerian signal (Figure 7), which we interpret as a star-spot moving into or out of view following the 11.68 day rotation period of the star (Barnes, 2007). We find that dropping this data from the no-GP model leads to an eccentricity distribution that essentially matches that of our GP model. The no-GP model is not equipped to recognize the high correlation between these points, and so the burst of high cadence data distorts the orbit modelāproducing both a non-zero eccentricity, and an overconfident posterior.
6em
@@ 4.3.4 Possible impacts to previous RV models
Feng etĀ al. (2023) used a moving average (MA) model to account for temporal correlation in the noise, which can be thought of as an simplified GP.
The MA model described in that work uses the past data points to predict the subset point, with for the HARPS (pre) data. We hypothesize that the MA model may be affected in a similar way to the no-GP model when it encounters a burst of hundreds of data points over a short period of time. Earlier data points containing different realizations of the stellar noise may fall outside of this -point memory, which could limit the models ability to recognize the temporal correlation and result in a higher eccentricity. It is possible that the MA model might arrive a low eccentricity when these points are dropped, just like our no-GP model.
@@ 4.3.5 Comparison with astrometry-only models
It was noted by Feng etĀ al. (2023) that ā..astrometry-only analyses have also resulted in high-eccentricity solutionsā¦.ā While the cited works of Benedict etĀ al. (2006) and FritzĀ Benedict (2022) did arrive at moderate eccentricities, we should clarify that they did so using joint models of RV and astrometry which likewise did not include a GP or other model of the stellar activity. In fact, no previous works have published an astrometry-only analysis to the best of our knowledge. In contrast, our true astrometry-only model finds an eccentricity with a 75% highest density interval of [0, 0.44], peaking at zero. The astrometry data therefore can be compatible with either the low or moderate eccentricity results.
@@ 4.3.6 Conclusion: support for low eccentricity
We find that the GP models do not appear to be over-fitting a real eccentricity signal, and in fact are more predictive of held out data than the simpler model not including a GP. Based on the improved predictive performance of the GP model, we conclude that the current data supports a low eccentricity orbit, consistent with zero to the precision of currently available data, in agreement with Mawet etĀ al. (2019), Llop-Sayson etĀ al. (2021), and Roettenbacher etĀ al. (2022).
@@ 4.4 Direction of Orbital Rotation
We find that our models of individual astrometric data sources do not strongly constrain the direction of apparent orbital rotation in the plane of the sky to be clockwise or counterclockwise. For instance, the model of the Hipparcos and Gaia DR3 proper motions is evenly split between the two directions (inclination below/above ). The model of the Hipparcos IAD does suggest that either the planet is moving counterclockwise and is near-coplanar with the disk, or, that the planet is moving clockwise but misaligned with the disk. That said, the combination of all data sources, particularly the Hubble FGS and Gaia DR3-DR2 acceleration, supports counterclockwise motion fairly conclusively (see Figure 4). This result is inconsistent with that of Feng etĀ al. (2023) which is clockwise as indicated in their Figure 2.
6em
@@ 4.5 Coplanarity with outer debris disk
We now compare our complete orbit model to the outer debris disk model fit to ALMA data by Booth etĀ al. (2023). We find that the HGCA, Hubble FGS, and DR3-DR2 all support a planet that is co-planar with the outer debris disk to in both inclination and position angle of ascending node. Investigating our models, we find that the data point pulling the model slightly away from co-planar is the Gaia-Hipparcos position difference. We hypothesize that if the planet is indeed co-planar, then a very slight miscalibration between the Hipparcos and Gaia velocity reference frames (as corrected by the HGCA), or a slight inaccuracy in our model of the HGCA position difference would account for this slight tension. This data point is particularly sensitive to systematics because of the very high proper motion of the star surfacing non-linearities in the R.A.-Dec. coordinate system and variable light travel time to the system.
An alternative explanation might come from the disk model itself, which assumes a perfectly symmetric disk. If this assumption were violated by a sufficiently clumpy substructure, it is possible that the revised position angle of ascending node could shift a small amount (M. Booth, private communications).
Regardless, when we compare the inclination of the complete model with that of the outer disk from Booth etĀ al. (2023), we find the a mutual inclination 75% HDI of . This is nearly consistent with coplanarity but we cannot rule out a modest misalignment.
-1.2cm-0.45cm
@@ 4.6 Mass
This mass is substantially higher than the value of reported by Llop-Sayson etĀ al. (2021), but this disagreement is only due to the different inclination ranges we findāadjusting the inclination of their posterior down to would give a mass of as well.
The mass we find is also higher than the value of reported by Feng etĀ al. (2023), but in this case the disagreement comes partially from the different reported inclination () and partially from the reported non-zero eccentricity of 0.26.
Based on the revised mass, we calculated an expected contrast ratio of the planet versus star to be in L band, and in M band using Sonora Bobcat models (Marley etĀ al., 2021) and an age of 439 Myr (Barnes, 2007). That said, these are nominal values only as there is considerable uncertainty in the age of the system.
@@ 4.7 Predictions
We present predictions in Figure 9 of the planetās apparent location for every year from 2018 through to 2029, in addition to four significant imaging epochs. From the RV data alone, the planetās location is confined to a narrow ring of mas at opposition, which occurs in 2024-12-07 + yr. We the addition of any individual absolute astrometry dataset, the position angle of the planet at conjunction is confined to be South of the star in 2025, North in 2029, and so on. With our full model incorporating all absolute astrometry data sources, the position is confined to a region just mas wide near opposition. With the revised inclination, an imaging detection of the planet away from opposition may also be feasible, as we find that it never appears closer to the star than mas. This last constraint does rely on the complete model of all absolute astrometry, so conservative imaging programs may still wish to time their observations near opposition.
@@ 4.8 Expected Constraints from Relative Astrometry
With regards to the orbit and mass determination, the main benefit of an imaging detection will be an accurate measurement of the relative astrometry between the planet and star. An imaging detection near opposition, while the most technically feasible would not contribute significant constraints to the inclination and therefore mass but would serve to confirm the orbitās position angle of ascending node. To constrain inclination and mass, we recommend an imaging epoch somewhat away from opposition. The strongest constraints on the planetās mass would come from a relative astrometry measurement recorded close to a quarter period away from conjunction, i.e. when the planet is closest to the star. Such an endeavor would of course have to be balanced with the detectability at this closer separation.
6em
@@ 5 Conclusions
In this work we have presented an extensive reanalysis and cross-validation of all available radial velocity and absolute astrometry data. We determine the period of the planetās orbit to be yr ( AU). We independently confirm the period using the first complete orbit model of the system from only absolute astrometry data. In doing so, we rule out the possibility that the 7.3 year periodic signal could be a result of stellar activity cycles instead of a planet.
We find a much lower overall inclination than previous works (, ) that is likely close to coplanar with the outer debris disk (Booth etĀ al., 2023), with the 75% HDI of the mutual inclination between . The higher inclination we find supports a dynamical mass of . We also determine that the planetās orbit has a low eccentricity of consistent with a near-circular orbit.
We discussed disagreements in the literature regarding the planetās time of conjunction / opposition, inclination, and eccentricity, and presented arguments for why our updated results differ.
Finally, we presented detailed predictions for the planetās apparent location versus time, broken down by different subsets of data.
Our significantly revised mass and position predictions will aid future efforts to detect the planet through imaging, and point to Eridani bĀ being a remarkably close analog to our own planet Jupiter.
We thank T. Brandt for guidance in modeling the HGCA. We thank William Robertson for discussions regarding second order effects for high proper motion and barycentric radial velocity stars. We also thank Eric Agol for discussions regarding Celerite.jl and Gaussian processes. We thank Jessica Speedie for discussions regarding ALMA data. We thank Mark Booth for sharing posteriors from Booth etĀ al. (2023) in machine-readable format.
@@ Appendix A Additional Tables
The residuals we digitized from Figure 7 of Benedict etĀ al. (2006) are also presented in Tables 4 and 5. Based on that figure, we interpret the coordinate as a negative change in declination, while is a negative change in right ascension, times the cosine of declination.

### v2

@@ 4.3.2 The GP results are robust to the choice of data
While our main models used separate GPs for each instrument, we also tested a model with a single GP for all data. This setup, related to a strategy suggested in Blunt et al. (2023), shares temporal correlation information between instruments, instead of treating data from each instrument as uncorrelated. We find essentially the same low-eccentricity results from this model, indicating that the GPs do not appear to be overfitting the data from each instrument.
@@ 4.3.3 High cadence data are the source of the disagreement
The HARPS (pre) series contains a burst of high cadence data, with 430 out of all 530 points captured over a 2-night range in October, 2004. This dense data from HARPS visibly traces out a correlated signal with a timescale much higher than the Keplerian signal (Figure 7), which we interpret as a star-spot moving into or out of view following the 11.68 day rotation period of the star (Barnes, 2007). We find that dropping this data from the no-GP model leads to an eccentricity distribution that essentially matches that of our GP model. The no-GP model is not equipped to recognize the high correlation between these points, and so the burst of high cadence data distorts the orbit model—producing both a non-zero eccentricity, and an overconfident posterior.
@@ 4.3.4 Possible impacts to previous RV models
Feng et al. (2023) used a moving average (MA) model to account for temporal correlation in the noise, which can be thought of as an simplified GP.
The MA model described in that work uses the past data points to predict the subset point, with for the HARPS (pre) data. We hypothesize that the MA model may be affected in a similar way to the no-GP model when it encounters a burst of hundreds of data points over a short period of time. Earlier data points containing different realizations of the stellar noise may fall outside of this -point memory, which could limit the models ability to recognize the temporal correlation and result in a higher eccentricity. It is possible that the MA model might arrive a low eccentricity when these points are dropped, just like our no-GP model does.
@@ 4.3.5 Comparison with astrometry-only models
It was noted by Feng et al. (2023) that “..astrometry-only analyses have also resulted in high-eccentricity solutions….” While the cited works of Benedict et al. (2006) and Fritz Benedict (2022) did arrive at moderate eccentricities, we should clarify that they did so using joint models of RV and astrometry which likewise did not include a GP or other model of the stellar activity. In fact, no previous works have published an eccentricity value from an astrometry-only analysis to the best of our knowledge. In contrast, our true astrometry-only model finds an eccentricity with a 75% highest density interval of [0, 0.44], peaking at zero. The astrometry data therefore can be compatible with either the low or moderate eccentricity results.
@@ 4.3.6 Conclusion: support for low eccentricity
We find that the GP models do not appear to be over-fitting a real eccentricity signal, and in fact are more predictive of held out data than the simpler model not including a GP. Based on the improved predictive performance of the GP model, we conclude that the current data supports a low eccentricity orbit, consistent with zero to the precision of currently available data, in agreement with Mawet et al. (2019), Llop-Sayson et al. (2021), and Roettenbacher et al. (2022).
@@ 4.4 Direction of Orbital Rotation
We find that our models of individual astrometric data sources do not strongly constrain the direction of apparent orbital rotation in the plane of the sky to be clockwise or counterclockwise. For instance, the model of the Hipparcos and Gaia DR3 proper motions is evenly split between the two directions (inclination below/above ). The model of the Hipparcos IAD does suggest that either the planet is moving counterclockwise and is near-coplanar with the disk, or, that the planet is moving clockwise but misaligned with the disk. That said, the combination of all data sources, particularly the Hubble FGS and Gaia DR3-DR2 acceleration, supports counterclockwise motion fairly conclusively (see Figure 4). This result is inconsistent with that of Feng et al. (2023) which is clockwise as indicated in their Figure 2.
@@ 4.5 Coplanarity with outer debris disk
We now compare our complete orbit model to the outer debris disk model fit to ALMA data by Booth et al. (2023). Their posteriors indicate the disk has an inclination of , and position angle of . We find that the HGCA, Hubble FGS, and DR3-DR2 all support a planet that is co-planar with the outer debris disk. Comparing the inclination of the complete model with that of the outer disk from Booth et al. (2023), we find a mutual inclination 75% HDI of .
It is worth noting that the uniform and sine priors we adopted for position angle of ascending node and inclination combine to produce an effective prior on the mutual inclination that goes to zero probability as the mutual inclination approaches zero. Given this, the mutual inclination we find is highly consistent with planet b being co-planar with the debis disk. Via the Savage–Dickey density ratio, we calculate a Bayes factor of approximately 130 for a model with a misalignment of , versus unconstrained mutual inclination. This is strong support in favor of alignment between the disk and planet.
We finally note that this conclusion is impacted by our inclusion of the DR2 proper motions. Although we have accounted for the bulk rotation between the DR2 and DR3 reference frames, there were additional changes to the processing of bright and high proper-motion stars between Gaia DR2 and (e)DR3 (Lindegren et al., 2021). These changes may result in an unknown shift between DR2 and DR3 that is unrelated to any underlying physical motion. The release of Gaia epoch astometry in a future data release promises to resolve this remaining question, as it will remove the need to compare data between Gaia releases.
@@ 4.6 Mass
This mass is substantially higher than the value of reported by Llop-Sayson et al. (2021), but this disagreement is only due to the different inclination ranges we find—adjusting the inclination of their posterior down to would give a mass of as well.
The mass we find is also higher than the value of reported by Feng et al. (2023), but in this case the disagreement comes partially from the different reported inclination () and partially from the reported non-zero eccentricity of 0.26.
Based on the revised mass, we calculated an expected contrast ratio of the planet versus star to be in L band, and in M band using Sonora Bobcat models (Marley et al., 2021) and an age of 439 Myr (Barnes, 2007). That said, these are nominal values only as there is considerable uncertainty in the age of the system.
@@ 4.7 Predictions
We present predictions in Figure 9 of the planet’s apparent location for every year from 2018 through to 2029, in addition to four significant imaging epochs. From the RV data alone, the planet’s location is confined to a narrow ring of mas at opposition, which occurs in 2024-12-07 + yr. We the addition of any individual absolute astrometry dataset, the position angle of the planet at conjunction is confined to be South of the star in 2025, North in 2029, and so on. With our full model incorporating all absolute astrometry data sources, the position is confined to a region just mas wide near opposition. With the revised inclination, an imaging detection of the planet away from opposition may also be feasible, as we find that it never appears closer to the star than mas. This last constraint does rely on the complete model of all absolute astrometry, so conservative imaging programs may still wish to time their observations near opposition.
@@ 4.8 Expected Constraints from Relative Astrometry
With regards to the orbit and mass determination, the main benefit of an imaging detection will be an accurate measurement of the relative astrometry between the planet and star. While an imaging detection near opposition would be the most technically feasible, it would not contribute significant constraints to the inclination and therefore mass. It would nonetheless serve to confirm the orbit’s position angle of ascending node. To constrain inclination and mass, we recommend an imaging epoch somewhat away from opposition. The strongest constraints on the planet’s mass would come from a relative astrometry measurement recorded close to a quarter period away from conjunction, i.e. when the planet is closest to the star. Such an endeavor would of course have to be balanced with the detectability at this closer separation.
@@ 5 Conclusions
In this work we have presented an extensive reanalysis and cross-validation of all available radial velocity and absolute astrometry data. We determine the period of the planet’s orbit to be yr ( AU). We independently confirm the period using the first complete orbit model of the system from only absolute astrometry data. In doing so, we rule out the possibility that the 7.3 year periodic signal could be a result of stellar activity cycles instead of a planet.
We find a much lower overall inclination than previous works (, ) that is consistent with coplanarity with the outer debris disk (, as derived from ALMA observations by (Booth et al., 2023), with the 75% HDI of the mutual inclination between . The higher inclination we find supports a dynamical mass of . We also determine that the planet’s orbit has a low eccentricity of consistent with a near-circular orbit.
We discussed disagreements in the literature regarding the planet’s time of conjunction and opposition; inclination; and eccentricity, and presented explanations for why our updated results differ.
Finally, we presented detailed predictions for the planet’s apparent location versus time, broken down by different subsets of data.
Our significantly revised mass and position predictions will aid future efforts to detect the planet through imaging, and point to Eridani b being a remarkably close analog to our own planet Jupiter.
We thank T. Brandt for guidance in modeling the HGCA. We thank William Roberson for discussions regarding second order effects for high proper motion and barycentric radial velocity stars. We also thank Eric Agol for discussions regarding Celerite.jl and Gaussian processes. We thank Jessica Speedie for discussions regarding ALMA data. We thank Mark Booth for sharing posteriors from Booth et al. (2023) in machine-readable format. We thank the anonymous referee for their insightful comments which have improved this work.
@@ Appendix A Additional Tables
The residuals we digitized from Figure 7 of Benedict et al. (2006) are also presented in Tables 4 and 5. Based on that figure, we interpret the coordinate as a negative change in declination, while is a negative change in right ascension, times the cosine of declination.